OpenStudy (anonymous):

-3/8 divided by (-9/10)

jimthompson5910 (jim_thompson5910):

Hint: -3/8 divided by (-9/10) is the same as -3/8 times -10/9

jimthompson5910 (jim_thompson5910):

When dividing fractions, you flip the second fraction and you multiply

OpenStudy (anonymous):

and i would get 30/72 right?

jimthompson5910 (jim_thompson5910):

can that be reduced? if so, what can it be reduced to?

OpenStudy (anonymous):

idk how can it be?

jimthompson5910 (jim_thompson5910):

what is the GCF of 30 and 72?

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

you can go higher

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

8 is not a factor of 30

jimthompson5910 (jim_thompson5910):

but notice how 6 is a factor of 30 and 72

jimthompson5910 (jim_thompson5910):

it turns out that the GCF is 6

OpenStudy (anonymous):

12 and 6

jimthompson5910 (jim_thompson5910):

so 30/6 = 5 72/6 = 12 which means 30/72 reduces to 5/12

OpenStudy (anonymous):

how do i do this one? -5/6 didved by 1/2?

jimthompson5910 (jim_thompson5910):

-5/6 divided by 1/2 is the same as -5/6 times 2/1

OpenStudy (anonymous):

and that would be -10/6 right?

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

that is correct, which can be reduced to -5/3

jimthompson5910 (jim_thompson5910):

since -10/2 = -5 and 6/2 = 3